OpenStudy (anonymous):

2y = x + 3 5y = x - 7 What is the solution set of the given system? a){(-1/3, -4/3)} b){(1/3, 4/3)} c){(-29/3, -10/3)} d){(29/3, 10/3)}

OpenStudy (anonymous):

Solve 1st equation - 2nd equation. that gives -3y=10. So, y =-10/3. Then solve for x using that y value. Plug into any equation. Plugging into 1st equation, for example, gives you -20/3=x+3. So, x = -29/3. Answer is (-29/3, -10/3), meaning the answer is (c).